# Machine-generated text detection prevents language model collapse

### v4

@@ Abstract
As Large Language Models (LLMs) become increasingly prevalent, their generated outputs are proliferating across the web, risking a future where machine-generated content dilutes human-authored text. Since online data is the primary resource for LLM pre-training, subsequent models could be trained on an unknown portion of synthetic samples. This will lead to model collapse, a degenerative process whereby LLMs reinforce their own errors, and ultimately yield a declining performance. In this study, we investigate the impact of decoding strategy on model collapse, analysing the characteristics of text at each model generation, the similarity to human references, and the resulting model performance. Using the decoding strategies that lead to the most significant degradation, we evaluate model collapse in more realistic scenarios where the origin of the data (human or synthetic) is unknown. We train a machine-generated text detector and propose an importance sampling approach to alleviate model collapse. Our method is validated on two LLM variants (GPT-2 and SmolLM2) on the open-ended text generation task. We demonstrate that it can not only prevent model collapse but also improve performance when sufficient human-authored samples are present. We release our code at https://github.com/GeorgeDrayson/model_collapse.
@@ 1 Introduction
Large Language Models (LLMs) can generate high-quality, human-like language across a wide range of applications. A key factor that drives their capabilities is the vast amount of data used to train them, which is predominantly based on text published on the web Wenzek et al. (2020a). The extensive adoption of LLMs will inevitably result in an ever-increasing amount of synthetic (machine-generated) data that will co-exist alongside or even dominate over human-generated text Dohmatob et al. (2024b), especially within online ecosystems such as social media platforms, news websites, and digital encyclopedias. Hence, there are legitimate concerns as to the effect this will have on future generations of language models trained on a mixed set of human and synthetic corpora.
Several works have attempted to simulate this scenario by recursively training language models on LLM-generated output Shumailov et al. (2023); Briesch et al. (2023); Alemohammad et al. (2024a). They show that this recursive training loop affects model performance, reflected by an increased perplexity among other metrics. The outcome of this recursive training process is referred to as “model collapse” Shumailov et al. (2023), a degenerative process caused by training on synthetic data from previous generations, leading to compounded errors and models that have converged to a low variance distribution that is significantly dissimilar to the original model trained on human data. This leads to models with significant performance degradation Alemohammad et al. (2024a) and a drastic loss in diversity Briesch et al. (2023); Guo et al. (2024); Alemohammad et al. (2024a). Related literature has proposed some ways to mitigate against this, such as using synthetic samples as negative guidance in the image domain Alemohammad et al. (2024b), selecting synthetic samples based on high perplexity scores Feng et al. (2024), token-level editing Zhu et al. (2024) or filtering low-quality samples Zhang et al. (2024). However, in these prior works, the models are either trained on entirely synthetic data or the labels of the samples are known a priori which is not necessarily reflective of more realistic settings.
Our work explores the scenario where the origin of the samples is unknown (human or synthetic). We propose a method for preventing model collapse by a guided resampling of the training samples using a machine-generated text detector. Our method is motivated by prior work Bertrand et al. (2024); Alemohammad et al. (2024a) which highlighted that when the proportion of human data in the training set is at a sufficient level, model collapse can be prevented. In addition to the current task-focused metrics used to assess model collapse, we evaluate the quality of the data from each model generation and its semantic resemblance to the original human samples. In this context, we also explore the effect of decoding strategy on the extent of model collapse and look into the characteristics of the synthetic data that could be causing this.
@@
    we present experiments on two LLMs across a range of decoding strategies.
@@ 2 Prior work on model collapse
Model collapse is a degenerative process in which models recursively trained on generational data exhibit a drop in performance compared to a model trained on the original human distribution Shumailov et al. (2023). In the early stages of recursive training, information is lost at the tails of the distribution and eventually, the output distribution converges to a point estimate with very little variance, resulting in a model that cannot be restored back to the original generation trained on human data. This effect can also be viewed as a change to neural scaling laws, in which there reaches a point where training on additional synthetic samples does not improve model performance and learning plateaus Dohmatob et al. (2024b).
@@
It has been argued that the two causes for this behaviour are finite sampling error leading to information being lost at the tails of the distribution, and functional approximation error introducing non-zero likelihoods outside of the support of the original distribution Shumailov et al. (2023). Additionally, Dohmatob et al. (2024b) theorised that the choice of generation algorithm is another contributing factor to model collapse. However, this has not been empirically evaluated in the case of LLMs, where decoding strategies that truncate the probability could have a significant impact. Currently, model collapse in LLMs has been studied with a fixed decoding strategy and model degradation has been mostly assessed using task performance metrics such as perplexity Shumailov et al. (2024) and test loss Gerstgrasser et al. (2024). Interestingly, Guo et al. (2024) also looks into the diversity of the generated text. Hence, in our study, we have chosen to study model collapse across three perspectives: the quality of the generated text (including diversity and readability), its similarity to human text, and the model task performance.
@@
Recent studies have explored methods for mitigating model collapse. Bertrand et al. (2024) and Alemohammad et al. (2024a) show that when a high enough proportion of human data is added to the training samples at each iteration, model collapse in diffusion models can be avoided. In the computational linguistics domain, Gerstgrasser et al. (2024) showed that by accumulating all cross-generational data and combining it with the original human data, model collapse can be significantly mitigated. Feng et al. (2024) and Zhang et al. (2024) investigated model collapse in a purely synthetic setting. They proposed heuristic methods for selecting samples, achieving performance akin to training on the original human data. However, in these prior works, the models are trained on either entirely synthetic data or the true labels of the samples are known. In our work, we investigate how to avoid model collapse in a more realistic setting where the training data is mixed and the origin (human or synthetic) of the samples is unknown.
@@ 3 Methods
3.1 Language models and decoding strategies
Language models learn a probability distribution $p_{\theta}\!\left(\mathbf{x}\right)$ for a sequence of tokens $\mathbf{x}\!=\!\{x_{1},\dots,x_{|\mathbf{x}|}\}$, where $\theta$ denotes the model parameters. In this work, we study open-ended text generation, in which a sequence of $m$ tokens, $\mathbf{x}\!=\!\{x_{1},\dots,x_{m}\}$, is provided as context and the task is to generate a continuation, $\mathbf{\hat{x}}\!=\!\{\hat{x}_{1},\dots,\hat{x}_{c}\}$, from the model’s probability distribution, as in
Tokens are selected from the probability distribution at each step by following a decoding strategy, resulting in a text sample $\{\mathbf{x},\hat{\mathbf{x}}\}$.
There are two main categories of decoding strategies, deterministic and stochastic. The former is designed to maximise the joint probability of the generated sequence, e.g. by selecting the most probable token at each step (greedy decoding) or keeping track of multiple candidate text sequences and selecting the most probable (beam search). This reduces the likelihood of converging to locally optimal but globally suboptimal sequences. Stochastic methods, on the other hand, produce less repetitive and more human-like text Holtzman et al. (2020). The simplest stochastic method samples directly from the distribution $p_{\theta}$. A more nuanced approach, referred to as top-$k$ decoding Fan et al. (2018), samples from the $k$ most probable tokens to avoid text generation from the tail of $p_{\theta}$. Nucleus sampling dynamically truncates the vocabulary to the highest probability tokens by thresholding the cumulative probability mass with a parameter $\eta\!\in\![0,1]$ Holtzman et al. (2020). Alternatively, the probability mass can be skewed towards high-probability outcomes by deploying temperature, controlled by parameter $\tau\!\in\![0,1]$ Ackley et al. (1985).
@@ 3.2 Recursive LLM training
Similarly to Shumailov et al. (2024) and Dohmatob et al. (2024b), we simulate model collapse by training a language model recursively on its own generated output (entirely or partially, depending on our underlying hypothesis) for a fixed number of generations. This process is described in Algorithm 1. Recursive training commences by fine-tuning a pre-trained language model, $p_{\theta}$, using a dataset consisting of $n$ human-generated samples, $\mathcal{D}_{\text{H}}\!=\!\{\mathbf{x}_{s}\}_{s=1}^{n}$. This results in a model $p^{0}$, where ‘00’ denotes the stage of the entire process (generation). [^1] For enhanced notational clarity, we choose to drop parameter $\theta$ for the recursively produced LLMs. However, we clarify that $\theta$ is updated in each generation. We then use a set of $n$ context sequences, $\mathcal{X}\!=\!\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ (one for each sample in $\mathcal{D}_{\text{H}}$), to generate a set of continuation sequences, $\mathcal{\hat{X}}\!=\!\{\mathbf{\hat{x}}_{1},\dots,\mathbf{\hat{x}}_{n}\}$, where $\hat{\mathbf{x}}_{s}\!\sim\!p^{0}$ (see also section 3.1). The human-generated context together with the LLM-generated continuation sequences form a new synthetic dataset, $\mathcal{D}_{\text{S}}^{1}$ (here ‘$1$’ is used to denote that this dataset will be used to fine-tune a language model in the next generation).
Subsequently, successive rounds of recursive training are carried out. In each generation $i$ with $i\!\geq\!1$, the original language model $p_{\theta}$ is fine-tuned using (an ever changing) synthetic dataset $\mathcal{D}_{\text{S}}^{i}$ to obtain $p^{i}$. Then, we prompt $p^{i}$ with context sequences $\mathcal{X}$ to generate a new synthetic dataset $\mathcal{D}_{\text{S}}^{i\!+\!1}$ that will be used to fine-tune $p_{\theta}$ in generation $i\!+\!1$.
@@ 3.3 LLM performance
We evaluate model collapse by fine-tuning and testing models on the open-ended text generation task, emulating the setup proposed by Shumailov et al. (2024) and Dohmatob et al. (2024b) to facilitate comparison. We assess language model performance in terms of perplexity and evaluation accuracy. Perplexity measures how well the model predicts unseen text, with lower values indicating better performance. Accuracy, in this context, reflects the proportion of correctly predicted tokens, providing a direct measure of the model’s effectiveness in generating accurate language outputs.
@@ 3.4 Metrics for LLM text generation quality
Self-BLEU Zhu et al. (2018) evaluates the BLEU score Papineni et al. (2002) of each document compared to all other documents in the generation set, providing a metric for how repetitive the model is for different outputs. Due to the computational expense of evaluating Self-BLEU we constrain the calculation to a random sample of $1{,}000$ documents and evaluate up to an $n$-gram size of $4$. A lower Self-BLEU score indicates higher text diversity.
Readability is evaluated using the Flesch-Kincaid Reading-Ease score Flesch (1948), which evaluates how difficult a passage is to understand based on the number of words, sentences, and syllables. We implement the metric using the textstat package.222textstat Python package, textstat.org. Lower scores indicate more complex text that contains longer sentences and verbose language.
@@ 3.5 Machine-generated text detection
Machine-generated text detection methods can be divided into neural-based Hu et al. (2023); Bhattacharjee et al. (2023) and metric-based approaches Mitchell et al. (2023); Hans et al. (2024). The former use statistical features, often extracted from surrogate LLMs, to detect machine-generated text, whereas the latter are based on machine learning such as fine-tuning a small pre-trained language model with a binary classification head. Here we deploy a neural classifier due to reported state-of-the-art (SOTA) performance on relevant machine-generated text detection benchmarks Wang et al. (2024a); Li et al. (2024). We use the output logits as a form of classification confidence.
Our detector is based on an encoder-only transformer model with a sequence classification head that maps the CLS token representation to logits, $\mathbf{z}_{i}$, which are converted to pseudo-probabilities using a sigmoid function, $\sigma$. As LLM training is considerably resource intensive, any data filtering or sampling methods must be able to efficiently process large quantities of data with minimal computational overhead Wenzek et al. (2020b). With this in consideration, we evaluated the base variants of $3$ pre-trained language models with under $200$ million parameters: RoBERTa Goyal et al. (2020) and DeBERTav3 He et al. (2023) due to their SOTA performance in machine-generated text detection Li et al. (2024); Wang et al. (2024b) and ModernBERT Warner et al. (2024) as a more recent variant that has achieved superior performance on a range of benchmarks. The added advantage of ModernBERT is the large context window ($8{,}192$ tokens), and its superior computational speed and memory efficiency Warner et al. (2024). See Appendix A for more details. Despite their strong performance, as with all modern neural networks, the confidence estimates are poorly calibrated Guo et al. (2017), i.e. the probability estimates are not representative of the true likelihood. To mitigate overconfidence in the predictions, we applied label smoothing. Additionally, we used temperature scaling to further calibrate the model’s predictions. Given the logit vector $\mathbf{z}_{i}$, the new confidence prediction is $\sigma\left(\mathbf{z}_{i}/T\right)$ where $T$ is a learnable temperature parameter.
@@
SmolLM2 greedy $13.36$ $90.24$ $47.89$ $44.00$ $6.99$ $3.02$ $61.90$ $50.23$ $4.29$ $1.12$ $62.33$ $\mathbf{46.72}$
beam search $13.36$ $88.90$ $47.87$ $43.96$ $6.75$ $2.88$ $61.47$ $\mathbf{44.47}$ $3.86$ $1.09$ $62.38$ $44.47$
sampling $13.36$ $14.76$ $47.88$ $46.82$ $90.83$ $\mathbf{91.62}$ $37.69$ $33.33$ $91.28$ $\mathbf{83.82}$ $47.49$ $44.33$
temperature $13.36$ $23.70$ $47.86$ $46.38$ $82.96$ $24.37$ $46.21$ $61.64$ $91.48$ $21.66$ $52.78$ $64.59$
top-$k$ $13.36$ $19.01$ $47.88$ $46.41$ $82.57$ $57.35$ $48.62$ $61.18$ $95.48$ $68.41$ $55.45$ $62.82$
nucleus $13.36$ $21.80$ $47.88$ $46.55$ $87.38$ $47.36$ $42.43$ $60.85$ $94.40$ $67.37$ $51.14$ $63.71$

### v5

@@ Abstract
As Large Language Models (LLMs) become increasingly prevalent, their generated outputs are proliferating across the web, risking a future where machine-generated content dilutes human-authored text. Since online data is the primary resource for LLM pre-training, subsequent models could be trained on an unknown portion of synthetic samples. This will lead to model collapse, a degenerative process whereby LLMs reinforce their own errors, converge to a low variance output distribution, and ultimately yield a declining performance. In this study, we investigate the impact of decoding strategy on model collapse, analysing the text characteristics at each model generation, the similarity to human references, and the resulting model performance. Using the decoding strategies that lead to the most significant degradation, we evaluate model collapse in more realistic scenarios where the origin of the data (human or synthetic) is unknown. We train a machine-generated text detector and propose an importance sampling approach to alleviate model collapse. Our method is validated on two LLM variants (GPT-2 and SmolLM2), across a range of model sizes ($124$M to $1.7$B), on the open-ended text generation task. We demonstrate that it can not only prevent model collapse but also improve performance when sufficient human-authored samples are present. Source code: github.com/GeorgeDrayson/model_collapse.
@@ 1 Introduction
Large Language Models (LLMs) can generate high-quality, fluent language across a wide range of applications. A key factor that drives their capabilities is the vast amount of data used to train them, which is predominantly based on text published on the web Wenzek et al. (2020). The extensive adoption of LLMs will inevitably result in an ever-increasing amount of synthetic data that will co-exist alongside or even dominate human-generated text Dohmatob et al. (2024), especially within online ecosystems such as social media, news websites, and digital encyclopedias. Hence, there are legitimate concerns as to the effect this might have on future generations of language models trained on a mixed set of human and synthetic corpora.
While synthetic data has proven beneficial in controlled scenarios, such as instruction tuning Wang et al. (2023) and distillation Hsieh et al. (2023), these settings typically involve careful curation and limited reuse. In contrast, our focus is on the long-term effects of uncontrolled accumulation of synthetic content. Several works have attempted to simulate this scenario by recursively training language models on LLM-generated output Shumailov et al. (2023); Briesch et al. (2023); Alemohammad et al. (2024a). The outcome of this recursive training is referred to as “model collapse” Shumailov et al. (2023), a degenerative process caused by training on synthetic data from previous generations, leading to compounded errors and the convergence to a low variance output distribution. This has been shown to cause performance degradation Alemohammad et al. (2024a) and a drastic loss in diversity Briesch et al. (2023); Guo et al. (2024); Alemohammad et al. (2024a). However, an unexplored factor in this recursive training process is the decoding strategy used to generate the synthetic data. Decoding strategies alter the distribution of model outputs, which could impact how errors accumulate during recursive training.
This work investigates the impact of decoding strategies on model collapse and the characteristics of the data that could be causing this. Subsequently, we explore the scenario where the training data is mixed (human and synthetic) in an unknown proportion, akin to training on web-crawled data. We propose a method for preventing model collapse by a guided resampling of the training data using a machine-generated text detector. Our method is motivated by prior work Bertrand et al. (2024); Alemohammad et al. (2024a), which highlighted that when the proportion of human data in the training set is sufficient, model collapse can be prevented.
@@
    we present experiments on two LLM variants across a range of decoding strategies and parameter sizes.
@@ 2 Prior work on model collapse
Model collapse is a degenerative process in which models recursively trained on generational data exhibit a drop in performance compared to a model trained on the original human distribution Shumailov et al. (2023). In the early stages of model collapse, information is lost at the tails of the distribution and eventually, the output distribution converges to a point estimate with very little variance, resulting in a model that cannot be restored back to the original generation trained on human data. This effect can also be viewed as a change to neural scaling laws, in which there reaches a point where training on additional synthetic samples does not improve model performance and learning plateaus Dohmatob et al. (2024).
@@
It has been argued that the two causes for this behaviour are finite sampling error leading to information being lost at the tails of the distribution, and functional approximation error introducing non-zero likelihoods outside of the support of the original distribution Shumailov et al. (2023). Additionally, Dohmatob et al. (2024) theorised that the choice of generation algorithm is another contributing factor to model collapse. However, this has not been empirically evaluated in the case of LLMs, where decoding strategies that modify the output distribution could have a significant impact. Currently, model collapse in LLMs has been studied with a fixed decoding strategy and model degradation has been mostly assessed using task performance metrics such as perplexity Shumailov et al. (2024) and test loss Gerstgrasser et al. (2024). Interestingly, Guo et al. (2024) also evaluate the diversity of the generated text. In our study, we have chosen to study model collapse across three perspectives: the quality of the generated text (including diversity and readability), its similarity to human text, and the model task performance.
@@
Recent studies have explored methods for mitigating model collapse, such as using synthetic samples as negative guidance in the image domain Alemohammad et al. (2024b), pruning samples based on high perplexity Feng et al. (2024), token-level editing Zhu et al. (2024) or filtering low-quality samples Zhang et al. (2024). Bertrand et al. (2024) and Alemohammad et al. (2024a) show that when a high enough proportion of human data is added at each iteration, model collapse in diffusion models can be avoided. In the computational linguistics domain, Gerstgrasser et al. (2024) showed that by accumulating all cross-generational data and combining it with the original human data, model collapse can be mitigated. However, in these works, the models are trained on either entirely synthetic data or the true labels of the samples are known a priori. In our work, we investigate how to avoid model collapse in a more realistic setting where the training data is mixed and the origin (human or synthetic) of the samples is unknown.
3 Background
In this work, we study open-ended text generation, in which a token sequence, $\mathbf{x}\!=\!\{x_{1},\dots,x_{m}\}$, is provided as context to a language model and the task is to generate a continuation, $\mathbf{\hat{x}}\!=\!\{\hat{x}_{1},\dots,\hat{x}_{c}\}$, from the model’s probability distribution, $p_{\theta}\!\left(\hat{\mathbf{x}}\right)$, where $\theta$ denotes the model’s parameters:
Tokens are selected from the probability distribution at each step by following a decoding strategy, resulting in a text sample $\{\mathbf{x},\hat{\mathbf{x}}\}$. There are two main categories of decoding strategies, deterministic and stochastic. The former is designed to maximise the joint probability of the generated sequence, e.g. by selecting the most probable token at each step (greedy decoding) or keeping track of multiple candidate text sequences and selecting the most probable (beam search). Stochastic methods, on the other hand, produce less repetitive and more human-like text Holtzman et al. (2020). The simplest stochastic method, pure sampling, samples directly from the distribution $p_{\theta}$. Top-$k$ decoding Fan et al. (2018), samples from the $k$ most probable tokens to avoid text generation from the tail of $p_{\theta}$. A more nuanced approach, nucleus sampling Holtzman et al. (2020), dynamically truncates the vocabulary to the highest probability tokens by thresholding the cumulative probability mass with a parameter $\eta\!\in\![0,1]$. Alternatively, the probability mass can be skewed towards high-probability outcomes by deploying temperature, controlled by $\tau\!\in\![0,1]$ Ackley et al. (1985).
@@ 4.1 Recursive LLM training
Similarly to Shumailov et al. (2024) and Dohmatob et al. (2024), we simulate model collapse by fine-tuning a language model recursively on its own generated output (entirely or partially, depending on our underlying hypothesis) for a fixed number of generations. This process is described in Algorithm 1. Recursive training commences by fine-tuning a pre-trained language model, $p_{\theta}$, using a dataset consisting of $n$ human-generated samples, $\mathcal{D}_{\text{H}}\!=\!\{\mathbf{x}_{s}\}_{s=1}^{n}$. This results in a model $p^{0}$, where ‘00’ denotes the stage of the entire process (generation). [^1] For enhanced notational clarity, we choose to drop parameter $\theta$ for the recursively produced LLMs. However, we clarify that $\theta$ is updated in each generation. We then use a set of $n$ context sequences, $\mathcal{X}\!=\!\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ (one for each sample in $\mathcal{D}_{\text{H}}$), to generate a set of continuation sequences, $\mathcal{\hat{X}}\!=\!\{\mathbf{\hat{x}}_{1},\dots,\mathbf{\hat{x}}_{n}\}$, where $\hat{\mathbf{x}}_{s}\!\sim\!p^{0}$ (see also section 3). The human-generated context together with the LLM-generated continuation sequences form a new synthetic dataset, $\mathcal{D}_{\text{S}}^{1}$ (here ‘$1$’ is used to denote that this dataset will be used to fine-tune a language model in the next generation).
Subsequently, successive rounds of recursive training are carried out. In each generation $i$, the original language model $p_{\theta}$ is fine-tuned using synthetic dataset $\mathcal{D}_{\text{S}}^{i}$ to obtain $p^{i}$. Thereafter, $p^{i}$ is prompted with context sequences $\mathcal{X}$ to generate a new synthetic dataset $\mathcal{D}_{\text{S}}^{i\!+\!1}$ that will be used to fine-tune $p_{\theta}$ in generation $i\!+\!1$.
@@ 4.2 LLM performance
We evaluate model collapse by fine-tuning and testing models on the open-ended text generation task, emulating the setup proposed by Shumailov et al. (2024) and Dohmatob et al. (2024). We assess language model performance in terms of perplexity and evaluation accuracy. Perplexity measures how well the model predicts unseen text, with lower values indicating better performance. Accuracy, in this context, reflects the proportion of correctly predicted tokens, providing a direct measure of the model’s effectiveness in generating accurate language outputs.
@@ 4.3 Metrics for LLM text generation quality
Self-BLEU Zhu et al. (2018) evaluates the BLEU score Papineni et al. (2002) of each document compared to all other documents in the generation set, providing a metric for how repetitive the model is for different outputs. We use a random sample of $1{,}000$ documents and evaluate Self-BLEU up to an $n$-gram size of $4$. A lower score indicates higher text diversity.
@@
Readability is evaluated using the Flesch-Kincaid Reading-Ease score Flesch (1948), which estimates how difficult it is to understand a passage based on the number of words, sentences, and syllables. We implement the metric using the textstat package.222textstat Python package, textstat.org. Lower scores indicate more complex text, typically characterised by longer sentences and higher lexical density.
@@ 4.4 Machine-generated text detection
Machine-generated text detection methods can be divided into neural-based Hu et al. (2023); Bhattacharjee et al. (2023) and metric-based approaches Mitchell et al. (2023); Hans et al. (2024). The former use statistical features, often extracted from surrogate LLMs, to detect machine-generated text, whereas the latter are based on machine learning, such as fine-tuning a small pre-trained language model with a binary classification head. Here we deploy a neural classifier due to reported state-of-the-art (SOTA) performance on relevant machine-generated text detection benchmarks Wang et al. (2024a); Li et al. (2024).
Our detector is based on an encoder-only transformer model with a sequence classification head that maps the CLS token representation to logits, $\mathbf{z}_{i}$, which are converted to pseudo-probabilities using a sigmoid function, $\sigma$. As LLM training is considerably resource-intensive, any data filtering or sampling methods must be able to efficiently process large quantities of data with minimal computational overhead Wenzek et al. (2020). With this in consideration, we evaluated the base variants of $3$ pre-trained language models with under $200$ million parameters: RoBERTa Goyal et al. (2020) and DeBERTav3 He et al. (2023) due to their SOTA performance in machine-generated text detection Li et al. (2024); Wang et al. (2024b) and ModernBERT Warner et al. (2024) as a more recent variant that has achieved superior performance on a range of benchmarks. The added advantages of ModernBERT is the large context window ($8{,}192$ tokens), superior computational speed, and memory efficiency Warner et al. (2024). See Appendix B for more details. Despite their strong performance, as with all modern neural networks, the confidence estimates are poorly calibrated Guo et al. (2017), i.e. they are not representative of the true likelihood. To mitigate overconfidence in the predictions, we applied label smoothing. Additionally, we used temperature scaling to further calibrate the model’s predictions. Given the logit vector $\mathbf{z}_{i}$, the new confidence prediction is $\sigma\!\left(\mathbf{z}_{i}/T\right)$ where $T$ is a learnable temperature parameter.
@@
SmolLM2 greedy $13.96$ $85.69$ $47.58$ $43.76$ $6.68$ $2.22$ $57.54$ $50.12$ $3.23$ $0.99$ $62.98$ $47.04$
beam search $13.96$ $84.75$ $47.59$ $43.77$ $6.64$ $2.16$ $57.30$ $50.38$ $3.06$ $0.89$ $62.69$ $45.87$
pure sampling $13.96$ $15.39$ $47.58$ $46.59$ $90.74$ $\mathbf{90.72}$ $47.23$ $\mathbf{45.66}$ $86.00$ $\mathbf{82.80}$ $47.55$ $\mathbf{47.59}$
temperature $13.96$ $24.88$ $47.55$ $46.05$ $82.92$ $24.81$ $51.15$ $57.55$ $89.94$ $17.60$ $52.83$ $64.09$
top-$k$ $13.96$ $19.86$ $47.59$ $46.02$ $82.72$ $56.77$ $52.62$ $57.91$ $85.97$ $59.84$ $55.52$ $63.18$
nucleus $13.96$ $22.81$ $47.58$ $46.23$ $87.33$ $44.27$ $49.62$ $58.39$ $92.39$ $53.00$ $51.20$ $64.17$